Given that \\(R=sin ¡t i+cos ¡t j+tk\\), find \\( (d^2 R)/(dt^2 )\\).
1
Expert's answer
2020-11-17T17:45:07-0500

First of all let us notice the fact that the basis vectors of "\mathbb{R}^3 (i, j, k)" are constant (as opposed to local vectors "e_r,e_\theta, e_z" of the cylindrical coordinates base, for example). Therefore the derivation of "R(t)" comes down to deriving every coordinate. Secondly, let us write "R(t)" as a vector for more obviousness:

"R(t)=\begin{pmatrix}\sin(\hat{a}t) \\ \cos(\hat{a}t) \\ t\end{pmatrix}"

Now let's calculate the first and then the second derivative:

"\frac{d}{dt} R(t) = \begin{pmatrix} \hat{a}\cos(\hat{a}t) \\ -\hat{a}\sin(\hat{a}t) \\ 1 \end{pmatrix}"

"\frac{d^2}{dt^2} R(t) = \begin{pmatrix} -\hat{a}^2 \sin(\hat{a}t) \\ -\hat{a}^2 \cos(\hat{a}t) \\ 0 \end{pmatrix}"